OpenStudy (anonymous):

I need help with a algebra 1 math question, its really confusing. someone please help quick

OpenStudy (cggurumanjunath):

post ur q !

OpenStudy (anonymous):

if you don't mind I have a couple questions but the first one is. ....

OpenStudy (cggurumanjunath):

ok.

OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 -4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16

OpenStudy (anonymous):

algebra is very hard for me I need a tutor lol

OpenStudy (cggurumanjunath):

ok. np. :)

OpenStudy (cggurumanjunath):

4x-2y=6 * /2 2x-y=3

OpenStudy (cggurumanjunath):

2x+y=5

OpenStudy (cggurumanjunath):

add 2 equations !

OpenStudy (anonymous):

im confused

OpenStudy (cggurumanjunath):

4x=8

OpenStudy (anonymous):

the question is up there and the answers a b c d and under them.

OpenStudy (cggurumanjunath):

4x-2y=6 2x+y=5 ..... these are the given equations

OpenStudy (cggurumanjunath):

on solving them as shown above we get x=2

OpenStudy (anonymous):

what you gave me doesn't fit any of my answers though

OpenStudy (cggurumanjunath):

d !

OpenStudy (anonymous):

4x + 2y = 10 8x = 16b ?

OpenStudy (cggurumanjunath):

2x-y=3

OpenStudy (cggurumanjunath):

2(2)-y=3

OpenStudy (cggurumanjunath):

4-3=1=y

OpenStudy (anonymous):

Aswer d is " 4x + 2y = 10 8x = 16 "

OpenStudy (cggurumanjunath):

yes .u r right !

OpenStudy (cggurumanjunath):

8x=16 * /2 x=2

OpenStudy (cggurumanjunath):

4x+2y=10 4(2)+2y=10 8+2y=10 2y=2 y=1

OpenStudy (anonymous):

My nxt question is " Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3) "

OpenStudy (cggurumanjunath):

so ,d is right !

OpenStudy (cggurumanjunath):

is this clear ?

OpenStudy (anonymous):

did you get my next question

OpenStudy (cggurumanjunath):

first tell me ! was the first q

OpenStudy (cggurumanjunath):

clear or not ?